Integrate dx/(25x²-4)½
1
Expert's answer
2020-02-20T09:48:29-0500

The integral is given by:

"\int \frac{dx}{\sqrt{25x^2 - 4}} \\"


Let us assume that:

"u = \frac{5x}{2}" (Equation 1)

"\Rightarrow du = \frac{5}{2} dx \\
\Rightarrow dx = \frac{2}{5} du \\"


From the (Equation 1), we can also write:

"x = \frac{2u}{5}"


Now:

Substituting "x = \frac{2u}{5}" into the integration and integrating with respect to "du" , we have:

"= \frac{2}{5} \int \frac{du}{\sqrt{4u^2 - 4}} \\
= \frac{2}{5} \int \frac{du}{2 \sqrt{u^2 - 1}} \\
= \frac{1}{5} \int \frac{du}{\sqrt{u^2 - 1}} \\"


Again:

Let us assume that:

"u = \sec (\theta ) \\
\Rightarrow du = \sec (\theta) \tan (\theta) \, d \theta \\"


Again:

Substituting "u = \sec (\theta)" into the integration and integrating with respect to "d \theta" , we have:

"= \frac{1}{5} \int \frac{\sec (\theta) \tan (\theta )}{\sqrt{\sec^2 (\theta) - 1}} \, d \theta \\"

"= \frac{1}{5} \int \frac{\sec (\theta) \tan (\theta)}{\tan (\theta)} \, d \theta" [ "\because \sec^2 (\theta) - \tan^2 (\theta) = 1" ]

"= \frac{1}{5} \int sec (\theta) \, d \theta \\
= \frac{1}{5} \ln \mid \sec (\theta) + \tan (\theta) \mid + C"

[ "\because \int \sec (x) \, dx = \ln \mid \tan (x) + \sec (x) \mid + C" ( Where C is a constant of integration) ]


Undo substitution "\sec (\theta ) = u" and "\tan (\theta) = \sqrt{u^2 -1}" into the integration and we have:

"= \frac{1}{5} \ln \mid u + \sqrt{u^2 - 1} \mid + C"


Again:

Undo substitution "u = \frac{5x}{2}" into the integration and we have:

"= \frac{1}{5} \ln \mid \frac{5x}{2} + \sqrt{ \frac{25x^2}{4} - 1 } \mid + C \\
= \frac{1}{5} \ln \mid \frac{5x}{2} + \frac{\sqrt{25x^2 -4}}{2} \mid + C \\"

"= \frac{1}{5} \ln \mid \frac{5x + \sqrt{25x^2 -4}}{2} \mid + C \\"

"= \frac{1}{5} \ln \mid 5x + \sqrt{25x^2 - 4} \mid - \frac{1}{5} \ln \mid 2 \mid + C" "\left[ \because \ln \mid \frac{a}{b} \mid = \ln \mid a \mid - \ln \mid b \mid \right]"

"= \frac{1}{5} \ln \mid 5x + \sqrt{25x^2 - 4} \mid + K" [ "\because K = C - \frac{1}{5} \ln \mid 2 \mid" ( Where K is an arbitrary constant) ]